OpenStudy (anonymous):

find the intersect point of y^2+x^2-16y+39=0 and y^2-x^2-9=0

OpenStudy (anonymous):

looks like you have a circle and a hyperbola right?

OpenStudy (anonymous):

maybe we can solve be solving both equations for \(x^2\) as a first step, although there might be an easier way

OpenStudy (anonymous):

ok well one is x^2=y^2-9 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes, this works out nicely how about the other one, you got that?

OpenStudy (anonymous):

yeah its x^2=-y^2+16y-39

OpenStudy (anonymous):

that is what i get to now we can solve \[y^2-9=-y^2+16y-39\] know how to do that?

OpenStudy (anonymous):

*too

OpenStudy (anonymous):

yeah one sec

OpenStudy (anonymous):

k suggestion: at one step, divide by 2 also it factors

OpenStudy (anonymous):

i did- y^2-9=-y^2+16y-39 +9 to both sides y^2=-y^2+16y-30 +y^2 to both sides 2y^2=16y-30 divide both sides by 2 y^2=8y-15 and thats as far as i understand what to do unless i did something wrong?

OpenStudy (anonymous):

looks good except that you have a quadratic equation (degree 2) so you want to set it equal to zero, i.e. write \[y^2-8y+15=0\] in order to solve it you can factor this one, so the zeros are not too hard to find i.e. you do not need the quadratic formula

OpenStudy (anonymous):

clear how to factor this one?

OpenStudy (anonymous):

im not very good with factoring

OpenStudy (anonymous):

think of two numbers whose product is \(15 \) and whose sum is \(-8\) not too many choices here

OpenStudy (anonymous):

if it is still not clear let me know and i will tell you

OpenStudy (anonymous):

3 and 5

OpenStudy (anonymous):

(y-3)(y-5)right?

OpenStudy (anonymous):

well, actually \(-3\) and \(-5\) fit the bill, since they have to add to \(-8\) but that means it factors as \[(x-3)(x-5)=0\] so the actual zero are \(3\) and \(5\)

OpenStudy (anonymous):

yes, you are right

OpenStudy (anonymous):

oops typo there, i means \(y\) not \(x\) sorry

OpenStudy (anonymous):

\[(y-3)(y-5)=0\implies y=3,y=5\] now we know two values for \(y\) but we still need to find \(x\)

OpenStudy (anonymous):

i think that i misinformed you in my question, im looking for the intersection points that the equations have together if not then keep going

OpenStudy (anonymous):

yeah i know that we are almost done

OpenStudy (anonymous):

oh ok cause you said zeros and i need the point that they intersect eachother lol

OpenStudy (anonymous):

i meant the zeros of the quadratic equation we had to solve we have two possible values for \(y\) namely \(3\) and \(5\) we can now find the corresponding values of \(x\) because we know \[x^2=y^2-9\]

OpenStudy (anonymous):

if \(y=3\) then this equation becomes \(x^2=3^2-9=0\) and so \(x^2=0\) tells us \(x=0\) one point of intersection is therefore \((0,3)\)

OpenStudy (anonymous):

now we have to be careful not to make a mistake with \(y=5\) get \[x^2=5^2-9=25-9=16\] i.e. \[x^2=16\]

OpenStudy (anonymous):

what is \(x\) ?

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

careful

OpenStudy (anonymous):

one possible answer is \(4\) and so one point of intersection is \((4,5)\) but there is another ...

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

\(x^2=16\) has two solutions, not just one

OpenStudy (anonymous):

you got one right away: \(x^2=16\) one answer is \(x=4\) you know the other one? (hint: it is negative)

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

right! remember i said be careful here we did not have to worry about it with the first one, because \(x^2=0\) has only one solution

OpenStudy (anonymous):

so there are 3 points of intersection: \[(-4,5), (0,3),(4,5)\]

OpenStudy (anonymous):

wanna check the answer?

OpenStudy (anonymous):

i can do it, THANK YOU FOR YOUR HELP!!!

OpenStudy (anonymous):

yw my pleasure i didn't mean actually check as in "plug in the numbers" i meant "cheat" http://www.wolframalpha.com/input/?i=y^2%2Bx^2-16y%2B39%3D0%2C+y^2-x^2-9%3D0

OpenStudy (anonymous):

solve the secold equation for x^2